Lucky while driving to school computes the average speed for his trip to be 20km/h on his return trip along the same route there is traffic and the average speed is 40km/h what is this average speed?
Answer by stanbon(75887)

Lucky while driving to school computes the average speed for his trip to be 20km/h on his return trip along the same route there is traffic and the average speed is 40km/h what is this average speed?
------
To school DATA:
rate = 20 kph ; distance = x km ; time = x/20 hrs
-----
From school DATA:
rate = 40 kph ; distance = x km ; time = x/40 hrs
-----
Average speed = total distance/total time = 2x/(x/20+x/40) = 2x/(60x/800)
----
= 1600/60 = 160/6 = 26 2/3 km/hr
Cheers,
Stan H.